how do you solve this equation 2(x+4)=2x-5?
Found 2 solutions by Edwin McCravy, ReadingBoosters:
Answer by Edwin McCravy(20077)

2(x+4) = 2x-5
2x+4 = 2x-5
Subtract 2x from both sides:
4 = -5
That's false so there is no solution. It's a contradictory equation.
Edwin

Answer by ReadingBoosters(3246)

Starting with
2(x+4)=2x-5
Distribute the 2
2x + 8 = 2x - 5
Combine like terms
2x - 2x = -5 - 8
0x = -13
x = -13/0: this expression is undefined
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com